Part A: A mixture of He, Ar, and Xe has a total pressure of 2.80 atm . The partial pressure of He is 0.250 atm , and the partial pressure of Ar is 0.450 atm . What is the partial pressure of Xe? Express your answer to three significant figures and include the appropriate units. Part B: A volume of 18.0 L contains a mixture of 0.250 mole N2 , 0.250 mole O2 , and an unknown quantity of He. The temperature of the mixture is 0 ∘C , and the total pressure is 1.00 atm . How many grams of helium are present in the gas mixture? Express your answer to three significant figures and include the appropriate units.
Answer #1

solution Payant- A Totod Pressure = 2.809/m. Partial Pressure of Heifhe) = 0.250 am Posetter Pressure of nr Pnv) = 0.450 cemmoles of He = Total mole of mixture -(mole of N₂ + mole of o, ) = 0.803 - (0- 250 +0.250) = 0.303 mole i molar mass of He = 4